# Regularity Propagation of Global Weak Solutions to a Navier–Stokes–Cahn–Hilliard System for Incompressible Two-phase Flows with Chemotaxis and Active Transport

### v1

@@ 3.2.2 Uniform estimates
$\leq C\|\nabla\partial_{t}\varphi^{\gamma,k}\|\|\partial_{t}\varphi^{\gamma,k}-\overline{\partial_{t}\varphi^{\gamma,k}}\|_{V_{0}^{\prime}}+C|\overline{\partial_{t}\varphi^{\gamma,k}}|^{2}$
$\leq\|\nabla\partial_{t}\varphi^{\gamma,k}\|\big{(}\|\nabla\mu^{\gamma,k}\|+\|\bm{v}^{k}\cdot\nabla\varphi^{\gamma,k}\|_{L^{\frac{6}{5}}}\big{)}+C|\overline{\partial_{t}\varphi^{\gamma,k}}|^{2}$
$\leq\frac{1}{20}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+C\|\nabla\mu^{\gamma,k}\|^{2}+C\|\bm{v}^{k}\|_{\bm{L}^{3}}^{2}\|\nabla\varphi^{\gamma,k}\|^{2}+C|\overline{\partial_{t}\varphi^{\gamma,k}}|^{2}$
$\leq\frac{1}{20}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+C\|\nabla\mu^{\gamma,k}\|^{2}+C\alpha^{2}|\overline{\varphi^{\gamma,k}}-c_{0}|^{2}+C\|\nabla\bm{v}^{k}\|^{2}.$
Thanks to the $L^{\infty}$-bound of $\varphi^{\gamma,k}$, we have
$I_{2}\leq\gamma\|\partial_{t}\bm{v}^{k}\|\|\nabla\partial_{t}\varphi^{\gamma,k}\|\|\varphi^{\gamma,k}\|_{L^{\infty}}\leq\frac{1}{20}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+\gamma^{2}C\|\partial_{t}\bm{v}^{k}\|^{2},$
@@
$I_{6}$ $\leq\theta_{0}\|\bm{v}^{k}\|\|\nabla\partial_{t}\varphi^{\gamma,k}\|\|\varphi^{\gamma,k}\|_{L^{\infty}}\leq\frac{1}{20}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+C\|\bm{v}^{k}\|^{2}.$
Next, exploiting the Gagliardo-Nirenberg inequality and the Sobolev embedding theorem, we infer from (3.39), (3.42), (3.47), (3.49) and (3.52) that
@@
$\leq\frac{1}{40}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+C\|\nabla\bm{v}^{k}\|^{2}\|\varphi^{\gamma,k}\|_{W^{2,4}}^{2}$
$\leq\frac{1}{40}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+C\|\nabla\bm{v}^{k}\|^{2}\big{(}1+\|\nabla\mu^{\gamma,k}\|^{2}+\|\sigma^{\gamma,k}\|_{L^{4}}^{2}+\gamma^{2}\|\partial_{t}\varphi^{\gamma,k}\|^{2}_{L^{4}}\big{)}$
$\leq\frac{1}{40}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+\gamma^{2}C\|\nabla\bm{v}^{k}\|^{2}\|\partial_{t}\varphi^{\gamma,k}\|^{\frac{1}{2}}\big{(}\|\nabla\partial_{t}\varphi^{\gamma,k}\|+|\overline{\partial_{t}\varphi^{\gamma,k}}|\big{)}^{\frac{3}{2}}$
$\quad+C\|\nabla\bm{v}^{k}\|^{2}\big{(}1+\|\nabla\mu^{\gamma,k}\|^{2}\big{)}$
$\leq\frac{1}{20}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+\gamma^{8}C\|\nabla\bm{v}^{k}\|^{8}\|\partial_{t}\varphi^{\gamma,k}\|^{2}+C\|\nabla\bm{v}^{k}\|^{2}\big{(}1+\|\nabla\mu^{\gamma,k}\|^{2}\big{)}$
$\quad+C\alpha^{2}|\overline{\varphi^{\gamma,k}}-c_{0}|^{2},$
@@
$\leq C\big{(}1+\|\nabla\mu^{\gamma,k}\|+\|\sigma^{\gamma,k}\|_{L^{3}}+\gamma\|\partial_{t}\varphi^{\gamma,k}\|_{L^{3}}\big{)}\|\nabla\bm{v}^{k}\|\|\nabla\partial_{t}\varphi^{\gamma,k}\|$
$\leq\frac{1}{40}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+\gamma^{2}C\|\partial_{t}\varphi^{k}\|\big{(}\|\nabla\partial_{t}\varphi^{\gamma,k}\|+|\overline{\partial_{t}\varphi^{\gamma,k}}|\big{)}\|\nabla\bm{v}^{k}\|^{2}$
$\quad+C\|\nabla\bm{v}^{k}\|^{2}\big{(}1+\|\nabla\mu^{\gamma,k}\|^{2}\big{)}$
$\leq\frac{1}{20}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+\gamma^{4}C\|\nabla\bm{v}^{k}\|^{4}\|\partial_{t}\varphi^{\gamma,k}\|^{2}+C\|\nabla\bm{v}^{k}\|^{2}\big{(}1+\|\nabla\mu^{\gamma,k}\|^{2}\big{)}$
$\quad+C\alpha^{2}|\overline{\varphi^{\gamma,k}}-c_{0}|^{2}.$
Using (3.40), (3.42), (3.52), (3.60) Hölder’s inequality and the Poincaré–Wirtinger inequality, we get
$I_{7}$ $=\chi\int_{\Omega}\big{(}\partial_{t}\varphi^{\gamma,k}-\overline{\partial_{t}\varphi^{\gamma,k}}\big{)}\partial_{t}\sigma^{\gamma,k}\,\mathrm{d}x$
$\leq C\|\partial_{t}\varphi^{\gamma,k}-\overline{\partial_{t}\varphi^{\gamma,k}}\|_{V_{0}}\|\partial_{t}\sigma^{\gamma,k}\|_{V_{0}^{\prime}}$
$\leq C\|\nabla\partial_{t}\varphi^{\gamma,k}\|\big{(}\|\nabla(\sigma^{\gamma,k}-\chi\varphi^{\gamma,k})\|+\|\bm{v}^{k}\|_{\bm{L}^{4}}\|\sigma^{\gamma,k}\|_{L^{4}}\big{)}$
$\leq\frac{1}{20}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+C\|\nabla(\sigma^{\gamma,k}-\chi\varphi^{\gamma,k})\|^{2}+C\|\nabla\bm{v}^{k}\|^{2}.$
In a similar manner, we can estimate $I_{8}$ as follows
@@
$\leq C\|\nabla\partial_{t}\varphi^{\gamma,k}\|\|\mathcal{N}(\partial_{t}\varphi^{\gamma,k}-\overline{\partial_{t}\varphi^{\gamma,k}})\|$
$\leq C\|\nabla\partial_{t}\varphi^{\gamma,k}\|\big{(}\|\bm{v}^{k}\cdot\nabla\varphi^{\gamma,k}\|_{L^{\frac{6}{5}}}+\|\mu^{\gamma,k}-\overline{\mu^{\gamma,k}}\|\big{)}$
$\leq\frac{1}{20}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+C\|\nabla\mu^{\gamma,k}\|^{2}+C\|\nabla\bm{v}^{k}\|^{2}.$
Recalling (3.42), (3.47), (3.48) and $\gamma\in(0,1)$, we can deduce that
@@
$\leq C\alpha^{2}|\overline{\varphi^{\gamma,k}}-c_{0}|\big{(}1+\|\nabla\mu^{\gamma,k}\|+\|\sigma^{\gamma,k}\|+\gamma\|\partial_{t}\varphi^{\gamma,k}\|\big{)}$
$\leq\|\nabla\mu^{\gamma,k}\|^{2}+\gamma\|\partial_{t}\varphi^{\gamma,k}\|^{2}+C\alpha^{2}|\overline{\varphi^{\gamma,k}}-c_{0}|+C\alpha^{4}|\overline{\varphi^{\gamma,k}}-c_{0}|^{2}$
By similar arguments for $I_{3}$ and $I_{7}$, we get
$I_{10}$ $\leq C\|\bm{v}^{k}\cdot\nabla\varphi^{\gamma,k}\|_{H^{1}}\|\partial_{t}\sigma^{\gamma,k}\|_{(H^{1})^{\prime}}$
$\leq C\|\nabla\bm{v}^{k}\|^{2}\|\nabla\varphi^{\gamma,k}\|_{\bm{L}^{\infty}}^{2}+C\|\bm{v}^{\gamma,k}\|_{\bm{L}^{6}}^{2}\|\varphi^{\gamma,k}\|_{W^{2,3}}^{2}+\|\partial_{t}\sigma^{\gamma,k}\|_{(H^{1})^{\prime}}^{2}$
$\leq C\|\nabla\bm{v}^{k}\|^{2}\|\varphi^{\gamma,k}\|_{W^{2,4}}^{2}+\|\partial_{t}\sigma^{\gamma,k}\|_{(H^{1})^{\prime}}^{2}$
$\leq\frac{1}{20}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+\gamma^{8}C\|\nabla\bm{v}^{k}\|^{8}\|\partial_{t}\varphi^{\gamma,k}\|^{2}+C\|\nabla\bm{v}^{k}\|^{2}\big{(}1+\|\nabla\mu^{\gamma,k}\|^{2}\big{)}$
$\quad+C\|\nabla(\sigma^{\gamma,k}-\chi\varphi^{\gamma,k})\|^{2}+C\alpha^{2}|\overline{\varphi^{\gamma,k}}-c_{0}|^{2},$
@@
$\leq\frac{1}{20}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+C\|\bm{v}^{k}\|^{2}.$
@@
$\mathcal{G}^{\gamma,k}(t)$ $=\frac{1}{2}\|\nabla\mu^{\gamma,k}(t)\|^{2}+\frac{\gamma}{2}\|\partial_{t}\varphi^{\gamma,k}(t)\|^{2}+\gamma\int_{\Omega}\bm{v}^{k}(t)\cdot\nabla\varphi^{\gamma,k}(t)\partial_{t}\varphi^{\gamma,k}(t)\,\mathrm{d}x$
@@
Thanks to (3.42), we have
$\left|\gamma\int_{\Omega}\bm{v}^{k}\cdot\nabla\varphi^{k}\partial_{t}\varphi^{k}\,\mathrm{d}x\right|$ $\leq\gamma\|\bm{v}^{k}\|_{\bm{L}^{\infty}}\|\nabla\varphi^{k}\|\|\partial_{t}\varphi^{k}\|$ (3.67)
$\leq\frac{\gamma}{8}\|\partial_{t}\varphi^{k}\|^{2}+\gamma C_{5}\|\bm{v}^{k}\|_{\bm{L}^{\infty}}^{2}.$
Noticing that $\gamma\in(0,1)$, a similar argument for $I_{9}$ yields
@@
$\quad\leq C\alpha|\overline{\varphi^{\gamma,k}}-c_{0}|\big{(}1+\|\nabla\mu^{\gamma,k}\|+\|\sigma^{\gamma,k}\|+\gamma\|\partial_{t}\varphi^{\gamma,k}\|\big{)}$
@@
$\quad\leq\frac{1}{4}\|\nabla\mu^{\gamma,k}\|^{2}+\frac{\gamma}{8}\|\partial_{t}\varphi^{\gamma,k}\|^{2}+C_{6}(\alpha+\alpha^{2})|\overline{\varphi^{\gamma,k}}-c_{0}|.$ (3.68)
@@
Combining the above estimates for $I_{1},\cdots,I_{11}$, we can deduce the following differential inequality from (3.66), (3.67) and (3.68):

### v2

@@ 3.2.2 Decoupled auxiliary problem for $(\varphi,\mu)$
$\leq C\|\partial_{t}\varphi^{\gamma,k}-\overline{\partial_{t}\varphi^{\gamma,k}}\|_{H^{1}}\|\partial_{t}\varphi^{\gamma,k}-\overline{\partial_{t}\varphi^{\gamma,k}}\|_{(H^{1})^{\prime}}+C|\overline{\partial_{t}\varphi^{\gamma,k}}|^{2}$
$\leq C\|\nabla\partial_{t}\varphi^{\gamma,k}\|\big(\|\nabla\mu^{\gamma,k}\|+\|\bm{v}^{k}\varphi^{\gamma,k}\|\big)+C|\overline{\partial_{t}\varphi^{\gamma,k}}|^{2}$
$\leq\frac{1}{18}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+C\|\nabla\mu^{\gamma,k}\|^{2}+C\|\bm{v}^{k}\|^{2}\|\varphi^{\gamma,k}\|_{L^{\infty}}^{2}+C|\overline{\partial_{t}\varphi^{\gamma,k}}|^{2}$
$\leq\frac{1}{18}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+C\|\nabla\mu^{\gamma,k}\|^{2}+C\alpha^{2}|\overline{\varphi^{\gamma,k}}-c_{0}|^{2}+C\|\bm{v}^{k}\|^{2}.$
Using the $L^{\infty}$-bound $\|\varphi^{\gamma,k}(t)\|_{L^{\infty}(0,T;L^{\infty}(\Omega))}\leq 1$ again, we have
$I_{2}\leq\gamma\|\partial_{t}\bm{v}^{k}\|\|\nabla\partial_{t}\varphi^{\gamma,k}\|\|\varphi^{\gamma,k}\|_{L^{\infty}}\leq\frac{1}{18}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+\gamma^{2}C\|\partial_{t}\bm{v}^{k}\|^{2},$
@@
$I_{6}$ $\leq\theta_{0}\|\bm{v}^{k}\|\|\nabla\partial_{t}\varphi^{\gamma,k}\|\|\varphi^{\gamma,k}\|_{L^{\infty}}\leq\frac{1}{18}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+C\|\bm{v}^{k}\|^{2}.$
Then exploiting the Gagliardo–Nirenberg inequality and the Sobolev embedding theorem, we infer from (3.68), (3.72), (3.73) and (3.74) that
@@
$\leq\frac{1}{36}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+C\|\nabla\bm{v}^{k}\|^{2}\|\varphi^{\gamma,k}\|_{W^{2,4}}^{2}$
$\leq\frac{1}{36}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+C\|\nabla\bm{v}^{k}\|^{2}\big(1+\|\nabla\mu^{\gamma,k}\|^{2}+\|\sigma^{k}\|_{L^{4}}^{2}+\gamma^{2}\|\partial_{t}\varphi^{\gamma,k}\|^{2}_{L^{4}}\big)$
$\leq\frac{1}{36}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+\gamma^{2}C\|\nabla\bm{v}^{k}\|^{2}\|\partial_{t}\varphi^{\gamma,k}\|^{\frac{1}{2}}\big(\|\nabla\partial_{t}\varphi^{\gamma,k}\|+|\overline{\partial_{t}\varphi^{\gamma,k}}|\big)^{\frac{3}{2}}$
$\quad+C\|\nabla\bm{v}^{k}\|^{2}\big(1+\|\nabla\mu^{\gamma,k}\|^{2}+\|\sigma^{k}\|_{L^{6}}^{2}\big)$
$\leq\frac{1}{18}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+\gamma^{8}C\|\nabla\bm{v}^{k}\|^{8}\|\partial_{t}\varphi^{\gamma,k}\|^{2}+C\|\nabla\bm{v}^{k}\|^{2}\|\nabla\mu^{\gamma,k}\|^{2}$
$\quad+C\|\nabla\bm{v}^{k}\|^{2}\big(1+\|\sigma^{k}\|_{L^{6}}^{2}\big)+C\alpha^{2}|\overline{\varphi^{\gamma,k}}-c_{0}|^{2},$
@@
$\leq C\big(1+\|\nabla\mu^{\gamma,k}\|+\|\sigma^{k}\|_{L^{3}}+\gamma\|\partial_{t}\varphi^{\gamma,k}\|_{L^{3}}\big)\|\nabla\bm{v}^{k}\|\|\nabla\partial_{t}\varphi^{\gamma,k}\|$
$\leq\frac{1}{36}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+\gamma^{2}C\|\partial_{t}\varphi^{\gamma,k}\|\big(\|\nabla\partial_{t}\varphi^{\gamma,k}\|+|\overline{\partial_{t}\varphi^{\gamma,k}}|\big)\|\nabla\bm{v}^{k}\|^{2}$
$\quad+C\|\nabla\bm{v}^{k}\|^{2}\big(1+\|\nabla\mu^{\gamma,k}\|^{2}+\|\sigma^{k}\|_{L^{6}}^{2}\big)$
$\leq\frac{1}{18}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+\gamma^{4}C\|\nabla\bm{v}^{k}\|^{4}\|\partial_{t}\varphi^{\gamma,k}\|^{2}+C\|\nabla\bm{v}^{k}\|^{2}\|\nabla\mu^{\gamma,k}\|^{2}$
$\quad+C\|\nabla\bm{v}^{k}\|^{2}\big(1+\|\sigma^{k}\|_{L^{6}}^{2}\big)+C\alpha^{2}|\overline{\varphi^{\gamma,k}}-c_{0}|^{2}.$
Next, using (3.68), (3.72), Hölder’s inequality and the Poincaré–Wirtinger inequality, we get
$I_{7}$ $=\chi\int_{\Omega}\big(\partial_{t}\varphi^{\gamma,k}-\overline{\partial_{t}\varphi^{\gamma,k}}\big)\partial_{t}\sigma^{k}\,\mathrm{d}x+\chi\overline{\partial_{t}\varphi^{\gamma,k}}\int_{\Omega}\partial_{t}\sigma^{k}\,\mathrm{d}x$
$\leq C\|\partial_{t}\varphi^{\gamma,k}-\overline{\partial_{t}\varphi^{\gamma,k}}\|_{H^{1}}\|\partial_{t}\sigma^{k}\|_{(H^{1})^{\prime}}+|\chi||\overline{\partial_{t}\varphi^{\gamma,k}}|\|1\|_{H^{1}}\|\partial_{t}\sigma^{k}\|_{(H^{1})^{\prime}}$
$\leq C\|\nabla\partial_{t}\varphi^{\gamma,k}\|\|\partial_{t}\sigma^{k}\|_{(H^{1})^{\prime}}+C\|\partial_{t}\sigma^{k}\|_{(H^{1})^{\prime}}$
$\leq\frac{1}{18}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+C(1+\|\partial_{t}\sigma^{k}\|_{(H^{1})^{\prime}}^{2}).$
In a similar manner, we can treat $I_{8}$ as follows
@@
$\leq|\beta|\|\partial_{t}\varphi^{\gamma,k}-\overline{\partial_{t}\varphi^{\gamma,k}}\|\|\mathcal{N}(\partial_{t}\varphi^{\gamma,k}-\overline{\partial_{t}\varphi^{\gamma,k}})\|$
$\leq C\|\nabla\partial_{t}\varphi^{\gamma,k}\|\big(\|\mu^{\gamma,k}-\overline{\mu^{\gamma,k}}\|+\|\bm{v}^{k}\cdot\nabla\varphi^{\gamma,k}\|_{L^{\frac{6}{5}}}\big)$
$\leq\frac{1}{18}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+C\|\nabla\mu^{\gamma,k}\|^{2}+C\|\nabla\bm{v}^{k}\|^{2}.$
Recalling (3.70), (3.72), (3.73), we estimate $I_{9}$ by
@@
$\leq C\alpha^{2}|\overline{\varphi^{\gamma,k}}-c_{0}|\big(1+\|\nabla\mu^{\gamma,k}\|+\|\sigma^{k}\|+\gamma\|\partial_{t}\varphi^{\gamma,k}\|\big)$
$\leq\|\nabla\mu^{\gamma,k}\|^{2}+\gamma^{2}\|\partial_{t}\varphi^{\gamma,k}\|^{2}+C(1+\|\sigma^{k}\|^{2})+C\alpha^{4}|\overline{\varphi^{\gamma,k}}-c_{0}|^{2}.$
Applying a similar argument for $I_{3}$ and $I_{7}$, we get
$I_{10}$ $\leq C\|\bm{v}^{k}\cdot\nabla\varphi^{\gamma,k}\|_{H^{1}}\|\partial_{t}\sigma^{k}\|_{(H^{1})^{\prime}}$
$\leq C\|\nabla\bm{v}^{k}\|^{2}\|\nabla\varphi^{\gamma,k}\|_{\bm{L}^{\infty}}^{2}+C\|\bm{v}^{k}\|_{\bm{L}^{6}}^{2}\|\varphi^{\gamma,k}\|_{W^{2,3}}^{2}+C\|\partial_{t}\sigma^{k}\|_{(H^{1})^{\prime}}^{2}$
$\leq C\|\nabla\bm{v}^{k}\|^{2}\|\varphi^{\gamma,k}\|_{W^{2,4}}^{2}+C\|\partial_{t}\sigma^{k}\|_{(H^{1})^{\prime}}^{2}$
$\leq\frac{1}{18}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+\gamma^{8}C\|\nabla\bm{v}^{k}\|^{8}\|\partial_{t}\varphi^{\gamma,k}\|^{2}+C\|\nabla\bm{v}^{k}\|^{2}\|\nabla\mu^{\gamma,k}\|^{2}$
$\quad+C\|\nabla\bm{v}^{k}\|^{2}\big(1+\|\sigma^{k}\|_{L^{6}}^{2}\big)+C\alpha^{2}|\overline{\varphi^{\gamma,k}}-c_{0}|^{2}+C\|\partial_{t}\sigma^{k}\|_{(H^{1})^{\prime}}^{2},$
@@
$\leq\frac{1}{18}\|\nabla\partial_{t}\varphi^{\gamma,k}\|^{2}+C\|\bm{v}^{k}\|^{2}.$
@@
$\mathcal{G}^{\gamma,k}(t)$ $=\frac{1}{2}\|\nabla\mu^{\gamma,k}(t)\|^{2}+\frac{\gamma}{2}\|\partial_{t}\varphi^{\gamma,k}(t)\|^{2}+\gamma\int_{\Omega}\big(\bm{v}^{k}(t)\cdot\nabla\varphi^{\gamma,k}(t)\big)\partial_{t}\varphi^{\gamma,k}(t)\,\mathrm{d}x$
@@
We need to control the last two terms in $\mathcal{G}^{\gamma,k}$ without a definite sign. It follows from (3.72) that
$\left|\gamma\int_{\Omega}(\bm{v}^{k}\cdot\nabla\varphi^{\gamma,k})\partial_{t}\varphi^{\gamma,k}\,\mathrm{d}x\right|$ $\leq\gamma\|\bm{v}^{k}\|_{\bm{L}^{\infty}}\|\nabla\varphi^{\gamma,k}\|\|\partial_{t}\varphi^{\gamma,k}\|$
$\leq\frac{\gamma}{8}\|\partial_{t}\varphi^{\gamma,k}\|^{2}+\gamma C_{12}\|\bm{v}^{k}\|_{\bm{L}^{\infty}}^{2}.$ (3.76)
On the other hand, for $\gamma\in(0,1)$, a similar argument to $I_{9}$ yields
@@
$\quad\leq C\alpha|\overline{\varphi^{\gamma,k}}-c_{0}|\big(1+\|\nabla\mu^{\gamma,k}\|+\|\sigma^{k}\|+\gamma\|\partial_{t}\varphi^{\gamma,k}\|\big)$
@@
$\quad\leq\frac{1}{4}\|\nabla\mu^{\gamma,k}\|^{2}+\frac{\gamma}{8}\|\partial_{t}\varphi^{\gamma,k}\|^{2}+C_{13}(1+\|\sigma^{k}\|^{2})+C_{13}\alpha^{2}|\overline{\varphi^{\gamma,k}}-c_{0}|^{2}.$ (3.77)
In (3.76), (3.77), the positive constants $C_{12}$, $C_{13}$ are independent of $\gamma$, $k$ and time. Recalling the construction of $\bm{v}^{k}$, $\sigma^{k}$, we choose the constants
$C_{14}(k,T)$ $=C_{12}\sup_{t\in[0,T]}\|\bm{v}^{k}(t)\|_{\bm{L}^{\infty}},$
$C_{15}$ $=C_{13}(1+\|\sigma\|_{L^{\infty}(0,+\infty;L^{2}(\Omega))}^{2})+4C_{13}\alpha^{2},$
so that
$\gamma C_{12}\|\bm{v}^{k}(t)\|_{\bm{L}^{\infty}}^{2}+C_{13}(1+\|\sigma^{k}(t)\|^{2})+C_{13}\alpha^{2}|\overline{\varphi^{\gamma,k}(t)}-c_{0}|^{2}\leq\gamma C_{14}(k,T)+C_{15},$
for all $t\in[0,T]$. We note that the constant $C_{14}\geq 0$ can depend on $k$ and $T$, while $C_{15}>0$ is independent of $\gamma$, $k$ and time. Define
$\widehat{\mathcal{G}^{\gamma,k}}(t)$ $=\mathcal{G}^{\gamma,k}(t)+\gamma C_{14}(k,T)+C_{15},$
which satisfies
$\widehat{\mathcal{G}^{\gamma,k}}(t)\geq\frac{1}{4}\|\nabla\mu^{\gamma,k}(t)\|^{2}+\frac{1}{4}\gamma\|\partial_{t}\varphi^{\gamma,k}(t)\|^{2},$
$\widehat{\mathcal{G}^{\gamma,k}}(t)\leq\frac{3}{4}\|\nabla\mu^{\gamma,k}(t)\|^{2}+\frac{3}{4}\gamma\|\partial_{t}\varphi^{\gamma,k}(t)\|^{2}+2\gamma C_{14}(k,T)+2C_{15},$
@@
for all $t\in[0,T]$. Combining the above estimates for $I_{1},\cdots,I_{11}$, we can deduce from (3.75) the following differential inequality: